# CNN-Enabled Scheduling for Probabilistic Real-Time Guarantees in Industrial URLLC

### v1

@@ Abstract
Index Terms:
Industrial wireless networks, URLLC, per packet real-time communications (PPRC), scheduling, CNN, convolutional neural networks, deep learning
@@ III-A CNN Architecture for Priority Prediction and Resource Block Selection Models
1:Input: $A_{i,1}$ (arrival time of first packet for link $i$), $M_{i}$ (set of interfering links of $i\in E$), $T_{l},D_{l}$ (period and relative deadline for $l\in M_{i}\cup\{i\}$), $X_{i,t}$ (local traffic demand at $i$), $\text{State}[l][\text{rb}][t]$ (transmission state for $l\in M_{i}\cup\{i\}$), $\text{Prio}[l][t]$ (priority for $l\in M_{i}\cup\{i\}$).
@@
3:$\text{State}[i][\text{rb}][t]=\text{UNDECIDED},\forall\text{rb}\in\text{RB}$
4:$\text{Prio}[i][t]=\text{CNN-Predict}(\{X_{i,t},d^{\prime\prime}_{i,t}-t,\text{State}[l][\text{rb}][t],\text{Prio}[l][t]\}_{l\in M_{i}})$
@@
[…]
@@
[…]
@@ IV-A System Evaluation and Performance Metrics
Schedulable $\text{Schedulable Ratio}=\dfrac{\sum_{i=1}^{N}\mathbb{1}\left(\frac{X_{i}}{C_{i}}\leq D_{i}\right)}{N}$
@@
[…]

### v3

@@ Abstract
publicationid: pubid: ©2025 IEEE . Personal use is permitted. Permission required for reuse.
@@ III-A CNN Architecture for Priority Prediction and Resource Block Selection Models
1:Input: $A_{i,1}$ (arrival time of first packet for link ii), $M_{i}$ (set of interfering links of $i\in E$), $T_{l},D_{l}$ (period and relative deadline for $l\in M_{i}\cup\{i\}$), $X_{i,t}$ (local traffic demand at ii), $\text{State}[l][\text{rb}][t]$ (transmission state for $l\in M_{i}\cup\{i\}$), $\text{Prio}[l][t]$ (priority for $l\in M_{i}\cup\{i\}$).
@@
[…]
@@
7:while $\text{done}==\text{false}$ do
8:  $\text{done}=\text{true}$
9:  for each $\text{rb}\in\text{RB}$ in increasing order of $\text{ID}(\text{rb})$ do
@@
19:   if $\text{State}[i][\text{rb}][t]==\text{UNDECIDED}$ and $\text{CNN-Rank}(\text{Prio}[i][t],\text{Prio}[l][t])$ holds for all $\text{UNDECIDED}\ l\in M_{i}$ then
@@ IV-A System Evaluation and Performance Metrics
$\text{Schedulable Ratio}=\dfrac{\sum_{i=1}^{N}\mathbb{1}\left(\frac{X_{i}}{C_{i}}\leq D_{i}\right)}{N}$
@@
The noise power $P_{\text{noise}}$ in a communication system is the power associated with the thermal noise in the system. It can be calculated using the following formula: $P_{\text{noise}}=kTB$ where: - $k$ is Boltzmann’s constant, $k=1.38\times 10^{-23}\,\text{J/K}$, - $T$ is the system temperature in Kelvin. A common value for communication networks is $T=290\,\text{K}$, which is considered room temperature, - $B$ is the bandwidth of the system in Hertz (Hz). We have used the same assumption as in [13] and assumed bandwidth $B=20\,\text{MHz}$.